# Avenida Diagonal Las Torres 2640, 7941169, Peñalolén, Chile

### v1

@@ 1 Introduction
Higher-curvature corrections in gravity theories appear in different contexts of theoretical physics. Their presence is inevitable if gravity is regarded as an effective theory from a Wilsonian viewpoint. The Gauss-Bonnet term, for instance, arises from the low-energy limit of string theory Zwiebach1985; Zwiebach2004. Higher-curvature corrections are also induced by quantum anomalies Graham:1999pm; AlvarezGaume:1983ig; Capper:1975ig; Deser:1993yx; Fernandes:1983ts; Gabadadze:2000kc and, in the case of quadratic gravity, the theory becomes one-loop renormalizable in Minkowski space, with the price of losing unitarity due to the presence of ghosts Stelle:1976gc. In the Anti-de Sitter/Conformal Field Theory (AdS/CFT) correspondence Maldacena:1997re; Aharony:1999ti, on the other hand, higher-curvature terms induce novel holographic sources at the conformal boundary, allowing one to compute quantum responses for the boundary theory Skenderis:2002wp. Also, the inclusion in the gravity action of terms of higher order in the curvature describes more generic CFT duals with different central charges, beyond what can be obtained in Einstein gravity Henningson:1998gx. However, the renormalization of infrared divergencies in asymptotically AdS spacetimes turns more challenging when higher-curvature terms are present. The resolution of this problem has not been achieved for generic gravitational solutions.
@@
Among the infinite family of theories with higher order in the curvature, Lovelock gravity Lovelock:1971yv can be regarded as the natural generalization of Einstein theory in higher dimensions. Its higher-curvature terms correspond to dimensionally-continued Euler topological densities and their field equations remain second-order, propagating the same number of degrees of freedom henneaux1990. The next term in the Lovelock series after the Ricci scalar is the Gauss-Bonnet density, which becomes dynamical in dimensions higher than four, generating interesting solutions such as black holes boulware1985; wheeler1986; birmingham1999; cvetic2002 and gravitational instantons Dehghani:2006aa; Corral:2019leh; Corral:2022udb; Corral:2025yvr, even with boundaries which are not conformally flat girao2014. On the other hand, the Gauss-Bonnet term contributes nontrivially to the black hole entropy, modifying the standard area law of Einstein gravity Jacobson:1993xs; Jacobson:1994MR; Wald:1993nt; Cai:2001dz.
@@
In the context of AdS/CFT, it has been shown that particular values of the Gauss-Bonnet coupling might violate the shear viscosity to entropy bound in the dual theory; a feature which has been conjectured to be universal in any holographic dual to Einstein-AdS gravity Brigante:2007nu. This provides a holographic argument for constraining the Gauss-Bonnet coupling such that the dual CFT is consistent. Additionally, in Refs. deBoer:2009pn; Buchel:2009sk the relation between the Gauss-Bonnet coupling and the universal coefficients arising in correlators of the holographic stress tensor in the dual CFT was studied. This implies a correction to the central charges, modifying the coefficients of the conformal anomaly in the dual CFT. Furthermore, it can be used to set causality and positivity bounds on the propagation of fluctuations as shown in Refs. deBoer:2009pn; Buchel:2009sk.
@@
Holographically, a renormalized gravitational on-shell action is needed for a well-defined generating functional of the dual CFT theory, in agreement with the GKPW formula Gubser:1998bc; Witten:1998qj. This requires the action to have a well-posed variational principle with respect to the holographic sources. This is also the case of Lovelock gravity in its linearized form.
@@
Motivated by the locality of the fields in the Wilsonian interpretation of QFT, the authors in Ref. Emparan:1999pm; hep-th/9910267; Balasubramanian:1999re; deHaro:2000vlm; Henningson:1998gx; Skenderis:2002wp, introduced the holographic renormalization prescription, which moreover ensures a Dirichlet variational principle with respect to the induced metric at finite radius $h_{ij}$. As a consequence, one has to add the Gibbons-Hawking-York (GHY) term in the case of Einstein gravity or the Myers terms Myers1987 for Lovelock gravity and then introduce a series of boundary covariant counterterms which cancel all the infrared divergences in a minimal subtraction scheme (i.e., without modifying the universal part).
@@
In the method outlined above, the equations of motion in the radial decomposition fix the subleading coefficients in the expansion (2) in terms of the holographic data $g_{(0)ij}$. In practice, however, finding the higher-order coefficients in terms of the $g_{(0)ij}$ using the field equations is technically hard for arbitrary higher-derivative gravity theories. Furthermore, the boundary terms required for fixing the Dirichlet variational principle with respect to the induced metric at finite radius are only known in the case of Lovelock gravity (Myers terms in Ref. Myers1987). For these reasons, the Holographic Renormalization (HR) counterterms are known for generic $g_{(0)ij}$ for Einstein-AdS in a low enough dimension. In turn, in Lovelock gravity, they are only known for a conformally flat boundary $g_{(0)ij}$ metric Yale:2011dq by assuming the form for the lowest-derivative counterterms, as covariant functionals of the boundary metric.
@@
For Einstein-AdS gravity in even bulk dimensions, a covariant renormalization scheme exists, which makes use of symmetry principles to ensure both the finiteness and the Dirichlet variational principle with respect to the $g_{(0)ij}$ source. This scheme is called Conformal Renormalization (CR) Anastasiou:2016jix; Anastasiou:2020mik; Anastasiou:2022ljq; Anastasiou:2023oro, and it considers embedding the theory into Conformal Gravity. Upon imposing the Einstein condition on the spacetime, the theory reduces to Einstein gravity with the necessary counterterms to achieve the renormalization. The method has the advantage that the counterterms are given directly from the embedding and are fixed. The justification for the prescription relies on the fact that all Einstein-AdS spacetimes are solutions of Conformal Gravity (CG) in four dimensions. In six dimensions, one can construct a specific combination of conformal invariants where this is still true. Then, as all Einstein solutions are also solutions of a given CG theory, one can choose this particular sector to evaluate the action. Then, the finiteness of Conformal Gravity for AAdS spaces gets inherited by the resulting corrected Einstein gravity action. The obtained action has a specific form, which considers the Einstein-Hilbert (EH) term, the topological Euler density with a specific coupling, and an extra boundary term that cancels remaining divergences that depend on the boundary Weyl tensor. The EH plus Euler part corresponds to the Topological Renormalization action Olea:2006vd; Miskovic:2009bm; 0708.0782; Miskovic:2014zja; Anastasiou:2020zwc, and the extra boundary term can be thought of as a correction, although it comes naturally from the CR prescription and therefore it is fixed by conformal symmetry. It is important to stress that the additional term is a fully covariant contribution proportional to $\Box\lvert F\rvert^{2}$, where
@@
The paper is organized as follows. Section 2 introduces Einstein-Gauss-Bonnet theory, discussing the equations of motion and vacua of the theory. The condition for the existence of a degenerate vacuum state is also analyzed. Section 2.2 considers the asymptotic analysis of the generic solutions of the theory, by introducing the FG expansion in terms of the holographic Poincaré coordinate z. The equations of motion are expanded in the FG gauge, and from them, the relation of the subleading coefficients of the metric in terms of the holographic data $g_{(0)ij}$ is determined (in the non-degenerate case). From this expansion, the bulk Weyl tensor is computed to all relevant orders in z, separating it into the AdS curvature $F$ defined in Eq. (3) and a tensor $X$ which is subleading when the two vacua are different. In Section 3, using the asymptotic form obtained for AAdS solutions to EGB, the Holographic Renormalization counterterms of the theory are worked out. Also, from the resulting renormalized EGB gravity action, the holographic stress-tensor is derived at the conformal boundary of the spacetime. Sections 4 and 5 present a different way of renormalizing the EGB-AdS theory in six dimensions, which gives identical results for holographic quantities at the conformal boundary but which is fully covariant in the spacetime metric. This procedure builds upon the addition of a topological term Olea:2005gb; Olea:2006vd; 0708.0782; Kofinas:2008ub; Miskovic:2009bm; Miskovic:2014zja; Anastasiou:2020zwc, but incorporates an extra total derivative which cancels remaining sub-leading divergences present for manifolds with non-conformally-flat boundaries. In Section 4, this total derivative is motivated by relating it to the conformal invariants in six dimensions, whereas the cancellation of the remaining divergences is verified in Section 5. The method has the advantage of facilitating the computation of holographic quantum information-theoretic measures Anastasiou:2017xjr; Anastasiou:2018mfk; Anastasiou:2018rla; Anastasiou:2019ldc; Anastasiou:2021swo; Anastasiou:2021jcv. It can be used for different metric gauges that are not of the FG form. Indeed, examples of non-Gaussian coordinates consider frames to study properties of conformal fluids in the context of the fluid/gravity correspondence Rangamani:2009xk. It also includes the possibility of an adapted gauge in the analysis of flat gravity, which has a well-defined asymptotically Minkowski limit Compere:2019bua. In Section 6, Boulware-Deser black holes in EGB theory with non-conformally flat transverse sections and non-constant curvature at the horizon are studied. These topological black holes are the generalization to EGB theory of the solutions for Einstein-AdS gravity analyzed in Ref. Anastasiou:2023oro. They serve as a test ground for computing their renormalized on-shell action using the scheme discussed in Section 5. Indeed, in Section 7, the conserved charges and thermodynamical quantities are computed for the black holes found in this work. Finally, Section 8 summarizes the main results and comments on their meaning and physical relevance.
@@ 2.1 Equations of motion
which gives a degenerate effective AdS radius with a value [^1] In five dimensions, in the so-called Chern-Simons point, the holographic description of the theory is radically different, what is reflected in the computation of the corresponding Weyl anomaly Banados:2005rz.
@@ 2.2 Asymptotic analysis
In order to work out the counterterms for EGB in AAdS spacetimes, one considers the FG expansion defined in Eqs. (1) and (2), using the effective AdS radius instead, what amounts to the change of $\ell\rightarrow\ell_{\rm eff}$. Then, this asymptotic expansion of the metric is replaced in the equations of motion (EOM) to find the subleading coefficients in terms of the boundary data $g_{(0)ij}$. Initially, the EOM are expressed in their independent components of the radial decomposition. To do so, one needs to introduce the Gauss-Codazzi relations, given in Appendix A. It is important to determine the expansion of the extrinsic curvature in this gauge
@@
For nondegenerate vacua, i.e., $\Delta^{\left(1\right)}\neq 0$, one can solve $g_{\left(2\right)ij}$ from the field equations explicitly. In such a case, it takes the form of Poincaré-Einstein manifolds Skenderis:2000in; deHaro:2000vlm, namely,
@@
The coefficient $g_{\left(2\right)ij}$ in Eq. (17) has a universal form (up to the definition of $\ell_{\rm eff}$), out of the degenerate point of the theory. In a similar fashion, from Eq. (19), one notices that $g_{\left(4\right)ij}$ can be decomposed into a universal part and a theory-dependent piece, with an explicit dependence on the Gauss-Bonnet coupling. The same conclusion can be drawn from a purely kinematic description, in terms to Penrose-Brown-Henneaux (PBH) transformations in the asymptotic form of the metric, mimicking the analysis in Ref. Brown:1986nw; Imbimbo:1999bj; Penrose:1985bww.
@@
There are features shared by both Einstein and EGB gravity, which become evident in the asymptotic behavior of the Weyl tensor. The latter can be decomposed on-shell as Jatkar:2015ffa
@@ 3.1 Dirichlet problem for the boundary metric
On the first place, the EGB action in Eq. (4) needs to be supplemented by a boundary term to determine a well-posed Dirichlet problem at constant radius. It can be considered as a generalization of the Gibbons-Hawking-York term for Lovelock theory. In particular, for the Gauss-Bonnet scalar, the Myers term Myers1987 can be cast as
@@
where $I_{\rm ct}$ indicates a series of covariant boundary terms that cancel the divergences asymptotically. These can be determined in a systematic way, as described for the first time in Ref. deHaro:2000vlm.
@@ 4 Conformal Invariants and Renormalization
whose coupling $\eta$ is fixed by requiring that the total action is zero for global AdS space Kofinas:2006hr
@@
In order to do so, the bulk action of Eq. (50) can be conveniently written as a polynomial in $F$ Anastasiou:2021jcv
@@
Notice that this action has a structure similar to that in Einstein-AdS6 gravity, renormalized by adding a topological term Miskovic:2014zja.
@@
In order to define a renormalized action for generic AAdS spaces, one requires additional boundary terms. In order to avoid any ad-hoc prescription, these counterterms are found by demanding bulk covariance and dependence on the Weyl tensor (as they should vanish for global AdS). Furthermore, they appear as a part of a conformal invariant in six dimensions. To start, a proper basis in that case is given by the invariants $\left\{I_{1},I_{2},I_{3}\right\}$ introduced in Ref. Bastianelli:2000rs, given by
@@
The reduced basis $\left\{\mathcal{I}_{1},\mathcal{I}_{2}\right\}$ of conformal invariants has a couple of notable features. On one hand, both invariants are expected to be finite on their own when evaluated in the FG frame. In particular, the finiteness of $\text{Pf}(W)$ for asymptotically hyperbolic spaces has been discussed in the mathematical literature Albin:2005qka. In the second invariant, the boundary term is responsible for the cancellation of the infrared divergences coming from the bulk. On the other hand, a linear combination $a_{1}\mathcal{I}_{1}+a_{2}\mathcal{I}_{2}$ for $a_{1}=\tfrac{1}{144}$ and $a_{2}=1$ corresponds to the conformal completion of Einstein-AdS action in six dimensions Anastasiou:2020mik; Anastasiou:2023oro. This is a Conformal Gravity theory with an Einstein sector, where the counterterms that appear are dictated by the conformal invariance of the action.

### v2

@@ 1 Introduction
Higher-curvature corrections in gravity theories appear in different contexts of theoretical physics. Their presence is inevitable if gravity is regarded as an effective theory from a Wilsonian viewpoint. The Gauss-Bonnet term, for instance, arises from the low-energy limit of string theory Zwiebach (1985, 2004). Higher-curvature corrections are also induced by quantum anomalies Graham and Witten (1999); Alvarez-Gaumé and Witten (1984); Capper and Duff (1975); Deser and Schwimmer (1993); Fernandes et al. (1983) and, in the case of quadratic gravity, the theory becomes one-loop renormalizable in Minkowski space, with the price of losing unitarity due to the presence of ghosts Stelle (1977). In the Anti-de Sitter/Conformal Field Theory (AdS/CFT) correspondence Maldacena (1999); Aharony et al. (2000), on the other hand, higher-curvature terms induce novel holographic sources at the conformal boundary, allowing one to compute quantum responses for the boundary theory Skenderis (2002). Also, the inclusion in the gravity action of terms of higher order in the curvature describes more generic CFT duals with different central charges, beyond what can be obtained in Einstein gravity Henningson and Skenderis (1998). However, the renormalization of infrared divergencies in asymptotically AdS spacetimes turns more challenging when higher-curvature terms are present. The resolution of this problem has not been achieved for generic gravitational solutions.
@@
Among the infinite family of theories with higher order in the curvature, Lovelock gravity Lovelock (1971) can be regarded as the natural generalization of Einstein theory in higher dimensions. Its higher-curvature terms correspond to dimensionally-continued Euler topological densities and their field equations remain second-order, propagating the same number of degrees of freedom Henneaux et al. (1990). The next term in the Lovelock series after the Ricci scalar is the Gauss-Bonnet density, which becomes dynamical in dimensions higher than four, generating interesting solutions such as black holes Boulware and Deser (1985); Wheeler (1986) and gravitational instantons Dehghani and Hendi (2006); Corral et al. (2019, 2022, 2025), even with boundaries which are not conformally flat. On the other hand, the Gauss-Bonnet term contributes nontrivially to the black hole entropy, modifying the standard area law of Einstein gravity Jacobson and Myers (1993); Wald (1993); Cai (2002).
@@
In the context of AdS/CFT, it has been shown that particular values of the Gauss-Bonnet coupling might violate the shear viscosity to entropy bound in the dual theory; a feature which has been conjectured to be universal in any holographic dual to Einstein-AdS gravity Brigante et al. (2008). This provides a holographic argument for constraining the Gauss-Bonnet coupling such that the dual CFT is consistent. Additionally, in Refs. de Boer et al. (2010); Buchel et al. (2010) the relation between the Gauss-Bonnet coupling and the universal coefficients arising in correlators of the holographic stress tensor in the dual CFT was studied. This implies a correction to the central charges, modifying the coefficients of the conformal anomaly in the dual CFT. Furthermore, it can be used to set causality and positivity bounds on the propagation of fluctuations as shown in Refs. de Boer et al. (2010); Buchel et al. (2010). [^1] Holographic aspects when the Gauss-Bonnet term is nonminimally coupled to scalar fields have been explored in Refs. Caceres et al. (2024); Tong et al. (2025).
@@
Holographically, a renormalized gravitational on-shell action is needed for a well-defined generating functional of the dual CFT theory, in agreement with the GKPW formula Gubser et al. (1998); Witten (1998). This requires the action to have a well-posed variational principle with respect to the holographic sources. This is also the case of Lovelock gravity in its linearized form.
@@
Motivated by the locality of the fields in the Wilsonian interpretation of QFT, the authors in Ref. Emparan et al. (1999); Imbimbo et al. (2000); Balasubramanian and Kraus (1999); de Haro et al. (2001); Henningson and Skenderis (1998); Skenderis (2002), introduced the holographic renormalization prescription, which moreover ensures a Dirichlet variational principle with respect to the induced metric at finite radius $h_{ij}$. As a consequence, one has to add the Gibbons-Hawking-York (GHY) term in the case of Einstein gravity or the Myers terms Myers (1987) for Lovelock gravity and then introduce a series of boundary covariant counterterms which cancel all the infrared divergences in a minimal subtraction scheme (i.e., without modifying the universal part).
@@
In the method outlined above, the equations of motion in the radial decomposition fix the subleading coefficients in the expansion (2) in terms of the holographic data $g_{(0)ij}$. In practice, however, finding the higher-order coefficients in terms of the $g_{(0)ij}$ using the field equations is technically hard for arbitrary higher-derivative gravity theories. Furthermore, the boundary terms required for fixing the Dirichlet variational principle with respect to the induced metric at finite radius are only known in the case of Lovelock gravity (Myers terms in Ref. Myers (1987)). For these reasons, the Holographic Renormalization (HR) counterterms are known for generic $g_{(0)ij}$ for Einstein-AdS in a low enough dimension. In turn, in Lovelock gravity, they are only known for a conformally flat boundary $g_{(0)ij}$ metric Yale (2011) by assuming the form for the lowest-derivative counterterms, as covariant functionals of the boundary metric.
@@
For Einstein-AdS gravity in even bulk dimensions, a covariant renormalization scheme exists, which makes use of symmetry principles to ensure both the finiteness and the Dirichlet variational principle with respect to the $g_{(0)ij}$ source. This scheme is called Conformal Renormalization (CR) Anastasiou and Olea (2016); Anastasiou et al. (2021c, 2022, 2023b), and it considers embedding the theory into Conformal Gravity. Upon imposing the Einstein condition on the spacetime, the theory reduces to Einstein gravity with the necessary counterterms to achieve the renormalization. The method has the advantage that the counterterms are given directly from the embedding and are fixed. The justification for the prescription relies on the fact that all Einstein-AdS spacetimes are solutions of Conformal Gravity (CG) in four dimensions. In six dimensions, one can construct a specific combination of conformal invariants where this is still true. Then, as all Einstein solutions are also solutions of a given CG theory, one can choose this particular sector to evaluate the action. Then, the finiteness of Conformal Gravity for AAdS spaces gets inherited by the resulting corrected Einstein gravity action. The obtained action has a specific form, which considers the Einstein-Hilbert (EH) term, the topological Euler density with a specific coupling, and an extra boundary term that cancels remaining divergences that depend on the boundary Weyl tensor. The EH plus Euler part corresponds to the Topological Renormalization action Aros et al. (2000b); Olea (2007); Miskovic and Olea (2009); Kofinas and Olea (2007); Miskovic et al. (2014); Anastasiou et al. (2020), and the extra boundary term can be thought of as a correction, although it comes naturally from the CR prescription and therefore it is fixed by conformal symmetry. It is important to stress that the additional term is a fully covariant contribution proportional to $\Box\lvert F\rvert^{2}$, where
@@
The paper is organized as follows. Section 2 introduces Einstein-Gauss-Bonnet theory, discussing the equations of motion and vacua of the theory. The condition for the existence of a degenerate vacuum state is also analyzed. Section 2.2 considers the asymptotic analysis of the generic solutions of the theory, by introducing the FG expansion in terms of the holographic Poincaré coordinate z. The equations of motion are expanded in the FG gauge, and from them, the relation of the subleading coefficients of the metric in terms of the holographic data $g_{(0)ij}$ is determined (in the non-degenerate case). From this expansion, the bulk Weyl tensor is computed to all relevant orders in z, separating it into the AdS curvature $F$ defined in Eq. (3) and a tensor $X$ which is subleading when the two vacua are different. In Section 3, using the asymptotic form obtained for AAdS solutions to EGB, the Holographic Renormalization counterterms of the theory are worked out. Also, from the resulting renormalized EGB gravity action, the holographic stress-tensor is derived at the conformal boundary of the spacetime. Sections 4 and 5 present a different way of renormalizing the EGB-AdS theory in six dimensions, which gives identical results for holographic quantities at the conformal boundary but which is fully covariant in the spacetime metric. This procedure builds upon the addition of a topological term Aros et al. (2000b, a); Olea (2005, 2007); Kofinas and Olea (2007, 2008); Miskovic and Olea (2009); Miskovic et al. (2014); Anastasiou et al. (2020), but incorporates an extra total derivative which cancels remaining sub-leading divergences present for manifolds with non-conformally-flat boundaries. In Section 4, this total derivative is motivated by relating it to the conformal invariants in six dimensions, whereas the cancellation of the remaining divergences is verified in Section 5. The method has the advantage of facilitating the computation of holographic quantum information-theoretic measures Anastasiou et al. (2018b, a, c, 2019, 2021b, 2021a). It can be used for different metric gauges that are not of the FG form. Indeed, examples of non-Gaussian coordinates consider frames to study properties of conformal fluids in the context of the fluid/gravity correspondence Rangamani (2009). It also includes the possibility of an adapted gauge in the analysis of flat gravity, which has a well-defined asymptotically Minkowski limit Compère et al. (2019). In Section 6, Boulware-Deser black holes in EGB theory with non-conformally flat transverse sections and non-constant curvature at the horizon are studied. These topological black holes are the generalization to EGB theory of the solutions for Einstein-AdS gravity analyzed in Ref. Anastasiou et al. (2023b). They serve as a test ground for computing their renormalized on-shell action using the scheme discussed in Section 5. Indeed, in Section 7, the conserved charges and thermodynamical quantities are computed for the black holes found in this work. Finally, Section 8 summarizes the main results and comments on their meaning and physical relevance.
@@ 2.1 Equations of motion
which gives a degenerate effective AdS radius with a value [^2] In five dimensions, in the so-called Chern-Simons point, the holographic description of the theory is radically different, which is reflected in the computation of the corresponding Weyl anomaly Banados et al. (2004, 2005).
@@ 2.2 Asymptotic analysis
In order to work out the counterterms for EGB in AAdS spacetimes, one considers the FG expansion defined in Eqs. (1) and (2), using the effective AdS radius instead, [^3] From a mathematical point of view, the FG expansion for Lovelock metrics was investigated in Refs. Albin (2009); Yu (2025). what amounts to the change of $\ell\rightarrow\ell_{\rm eff}$. Then, this asymptotic expansion of the metric is replaced in the equations of motion (EOM) to find the subleading coefficients in terms of the boundary data $g_{(0)ij}$. Initially, the EOM are expressed in their independent components of the radial decomposition. To do so, one needs to introduce the Gauss-Codazzi relations, given in Appendix A. It is important to determine the expansion of the extrinsic curvature in this gauge
@@
For nondegenerate vacua, i.e., $\Delta^{\left(1\right)}\neq 0$, one can solve $g_{\left(2\right)ij}$ from the field equations explicitly. In such a case, it takes the form of Poincaré-Einstein manifolds Skenderis (2001); de Haro et al. (2001), namely,
@@
The coefficient $g_{\left(2\right)ij}$ in Eq. (17) has a universal form (up to the definition of $\ell_{\rm eff}$), out of the degenerate point of the theory. In a similar fashion, from Eq. (19), one notices that $g_{\left(4\right)ij}$ can be decomposed into a universal part and a theory-dependent piece, with an explicit dependence on the Gauss-Bonnet coupling. The same conclusion can be drawn from a purely kinematic description, in terms to Penrose-Brown-Henneaux (PBH) transformations in the asymptotic form of the metric, mimicking the analysis in Ref. Brown and Henneaux (1986); Imbimbo et al. (2000); Penrose and Rindler (2011).
@@
There are features shared by both Einstein and EGB gravity, which become evident in the asymptotic behavior of the Weyl tensor. The latter can be decomposed on-shell as Jatkar et al. (2015)
@@ 3.1 Dirichlet problem for the boundary metric
On the first place, the EGB action in Eq. (4) needs to be supplemented by a boundary term to determine a well-posed Dirichlet problem at constant radius. It can be considered as a generalization of the Gibbons-Hawking-York term for Lovelock theory. In particular, for the Gauss-Bonnet scalar, the Myers term Myers (1987) can be cast as
@@
where $I_{\rm ct}$ indicates a series of covariant boundary terms that cancel the divergences asymptotically. These can be determined in a systematic way, as described for the first time in Ref. de Haro et al. (2001).
@@ 4 Conformal Invariants and Renormalization
whose coupling $\eta$ is fixed by requiring that the total action is zero for global AdS space Kofinas and Olea (2006)
@@
In order to do so, the bulk action of Eq. (50) can be conveniently written as a polynomial in $F$ Anastasiou et al. (2021a)
@@
Notice that this action has a structure similar to that in Einstein-AdS6 gravity, renormalized by adding a topological term Miskovic et al. (2014).
@@
In order to define a renormalized action for generic AAdS spaces, one requires additional boundary terms. In order to avoid any ad-hoc prescription, these counterterms are found by demanding bulk covariance and dependence on the Weyl tensor (as they should vanish for global AdS). Furthermore, they appear as a part of a conformal invariant in six dimensions. To start, a proper basis in that case is given by the invariants $\left\{I_{1},I_{2},I_{3}\right\}$ introduced in Ref. Bastianelli et al. (2001), given by
@@
The reduced basis $\left\{\mathcal{I}_{1},\mathcal{I}_{2}\right\}$ of conformal invariants has a couple of notable features. On one hand, both invariants are expected to be finite on their own when evaluated in the FG frame. In particular, the finiteness of $\text{Pf}(W)$ for asymptotically hyperbolic spaces has been discussed in the mathematical literature Albin (2009). In the second invariant, the boundary term is responsible for the cancellation of the infrared divergences coming from the bulk. On the other hand, a linear combination $a_{1}\mathcal{I}_{1}+a_{2}\mathcal{I}_{2}$ for $a_{1}=\tfrac{1}{144}$ and $a_{2}=1$ corresponds to the conformal completion of Einstein-AdS action in six dimensions Anastasiou et al. (2021c, 2023b). This is a Conformal Gravity theory with an Einstein sector, where the counterterms that appear are dictated by the conformal invariance of the action.